OpenStudy (anonymous):

Can sumone check my answer on this problem?: Solve: 2(x - 3) + 2(x - 7) = 40 A. x = 12.5 B. x = 15 C. x = 5 D. no solution E. infinitely many solutions

jhonyy9 (jhonyy9):

x=15

jhonyy9 (jhonyy9):

x=60/4 x=15

OpenStudy (anonymous):

\[2x+2x=14+6+20\] \[4x=60\] \[So, x=15\] \[Option, B\]

OpenStudy (anonymous):

thnk u guys

Directrix (directrix):

B. x = 15